OpenStudy (adamaero):

If f(x) varies directly with x^2, and f(x) = 72 when x = 6, find the value of f(3) MEDAL & maybe fan if you explain good =J

myininaya (myininaya):

\[f(x)=kx^2 \text{ by definition of } f(x) \text{ varies directly with } x^2 \\ 72=k(6)^2 \text{ solve for the constant k }\] put in the constant k in to f(x)=kx^2 then replace x with 3 to find f(3)

myininaya (myininaya):

But I don't believe this is your question, is it @adamaero ?

OpenStudy (adamaero):

I don't know. I re-posted this question, but then closed it because...let's just say it's a long story. I got it wrong for someone else. See: http://openstudy.com/study#/updates/5487782ce4b0337e71f8897c